OpenStudy (anonymous):

Please help! :) Find all real solutions : x/2x+7-x+1/x+3=1

OpenStudy (anonymous):

\[\frac{ x }{ 2x+7 }- \frac{ x+1 }{ x+3 }=1\]

OpenStudy (anonymous):

herro

OpenStudy (anonymous):

I would first find common denominator which is (2x+7)(x+3)

OpenStudy (anonymous):

ok i did that

OpenStudy (anonymous):

which would give us....x(x+3)-(2x+7)(x+1)=(2x+7)(x+3)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

hello